A fruit juice company mixes pineapple juice that sells for $5.50 per gallon with 100 gallons of orange juice that sells for $3.00 per gallon. How much pineapple juice is used to make a pineapple orange juice drink selling for $3.50 per gallon? Thank you
Answer by nyc_function(2741)

P = pineapple juice @ $5.50/gal
O = orange juice @ $3.00/gal
T = target @ $3.50/gal
▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬
the ratio of P & O needed will be in inverse proportion to distance from T
▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬▬
P:O = (3.50 - 3.00) : (5.50 - 3.50) = 0.50:2 = 25:100
ans:
▬▬
25 gal of pineapple juice
-----------------------------------
the conventional method is always there !
suppose x gal of pineapple juice, then
5.5x + 3*100 = 3.5*125
5.5x = 437.5 - 300 = 137.5
x = 137.5 /5.5
= 25 gal
------------
but surely more complex !